OpenStudy (anonymous):

A square pyramid has a slant height of 17 inches and a height of 15 inches. What is its surface area if the square base has sides of length 16 inches? 224 in2 480 in2 800 in2 864 in2

OpenStudy (anonymous):

If a square pyramid's base has a side 's' and its slant height is 'l', its total surface area is given by, s^2 + 2sl.................. see that its vertical height isn't used = 16^2 + 2 * 16 * 17 =256 + 544 =800 square inches

OpenStudy (anonymous):

thnx!!!!